OpenStudy (anonymous):

Change to a decimal: 14 1/2 %.

OpenStudy (anonymous):

0.145

OpenStudy (anonymous):

\[\frac{14+\frac{1}{2}}{100}=\frac{29}{200}=0.145 \]